# Perceptive Pedipulation with Local Obstacle Avoidance

### v1

@@ Abstract
Pedipulation leverages the feet of legged robots for mobile manipulation, eliminating the need for dedicated robotic arms. While previous works have showcased blind and task-specific pedipulation skills, they fail to account for static and dynamic obstacles in the environment. To address this limitation, we introduce a reinforcement learning-based approach to train a whole-body obstacle-aware policy that tracks foot position commands while simultaneously avoiding obstacles. Despite training the policy in only five different static scenarios in simulation, we show that it generalizes to unknown environments with different numbers and types of obstacles. We analyze the performance of our method through a set of simulation experiments and successfully deploy the learned policy on the ANYmal quadruped, demonstrating its capability to follow foot commands while navigating around static and dynamic obstacles.
@@ I Introduction
Recent works have explored pedipulation – manipulation using a quadrupedal robot’s foot – to overcome these limitations [5, 6, 7]. These works use reinforcement learning (RL) to track position commands for the pedipulating foot with a whole-body behavior. They successfully demonstrate skills such as pushing a button, opening a door, and lifting objects using the foot as an end-effector. However, these methods primarily rely on proprioceptive information and do not consider obstacles in the environment. Thus, the learned behaviors are unable to avoid unwanted collisions.
A potential approach to address this limitation is to develop a hierarchical architecture. In this setup, a learned blind pedipulation policy could serve as the low-level controller, while a human operator, classical path planner, or RL planner [8, 9, 10] could act as a high-level obstacle-aware planner, that provides collision-free trajectories for the foot end-effector. However, since the high-level planner can only command the foot positions and not control the entire system, collisions may still occur with the other parts of the robot during command tracking by the blind policy. Implementing large safety margins around the robot’s collision bodies can mitigate this issue, but it makes applications in confined spaces infeasible.
Thus, instead of a hierarchical design, this work proposes an end-to-end approach to learn a perceptive pedipulation policy. This policy retains the simple interface from prior works [5, 6, 7], requiring only a foot position command while extending the system’s capability to simultaneously perform whole-body obstacle avoidance. We train this policy entirely in simulation using simple arrangements of cuboidal obstacles and providing appropriate collision rewards to learn the desired collision avoidance behavior. In addition, we introduce a switch to allow or disallow contacts with the pedipulating foot. With this contact switch, the policy can either avoid obstacles entirely or perform pedipulation tasks while avoiding obstacles with the rest of the body, retaining the capabilities from our previous work [5].
@@ II Related Work
For general and precise manipulation tasks, recent works have looked into pedipulation - the manipulation of objects using a robot’s foot [6, 7, 5, 15]. These works specify a desired position for the pedipulating foot and control it with learning-based methods in a whole-body fashion. This approach provides a simple control interface for a human operator or a high-level planner to perform various tasks. However, the learned pedipulation policy is limited to obstacle-free environments, as it relies only on the proprioceptive information. In this work, we aim to account for obstacles in the robot’s workspace by incorporating perceptive observations.
@@
Recently, researchers are also investigating learning-based methods for obstacle avoidance [20]. Miki et al. [1] use perception to enable collision-free locomotion in confined spaces by training a high-level policy that commands the base pose and using a low-level policy to track this command. As motivated in Sec. I, this decoupling is limiting in the case of pedipulation. Honerkamp et al. [21] learn obstacle-avoiding manipulation skills for a wheeled-mobile base with an attached manipulator arm. In their work, the learned policy outputs the end-effector and base velocity commands, which are converted to the desired joint positions through inverse kinematics (IK). While it is possible to follow a similar approach that separates the system into a three-legged mobile base and the pedipulating leg, using IK reduces the whole-body reachability, as shown in [5].
@@ III-A Training Overview
We add Stage 3 to improve the sim-to-real transfer, as the elevation map on the robot is subject to noise, outliers, and occlusions.
@@ III-B1 Obstacles
Previous works integrate obstacles into the static terrain [10, 23], which is computationally efficient as it avoids the computation of the obstacle dynamics. In the context of pedipulation, however, where moving obstacles should be possible when desired (Sec. III-D), having only static obstacles would prevent the policy from learning interactive behaviors. Thus, we use movable obstacles, which increase training times but provide realistic collision dynamics.
@@ III-B2 Obstacle Scenarios
To encourage the policy to learn generalizable obstacle avoidance, we construct five obstacle scenarios (Fig. 2, Stage 2). We construct two scenarios using a single obstacle (Table LABEL:tab:obstacle_rand, Obstacle 1), focusing on close and far-range obstacle avoidance (Fig. 2-C and Fig. 2-D, respectively). Our third scenario (Fig. 2-E) has two obstacles (2 and 3) wide enough for the robot to spawn in between, which encourages the policy to act conservatively when constrained on both sides. Similarly, in the next environment (Fig. 2-F), we use two obstacles (2 and 3) to create a tight gap, into which the robot can only reach with its foot. Finally, using all three obstacles, we construct a scenario that features two gaps to reach through and position the obstacles such that there is enough space for the robot to turn on the spot to reach commands behind it (Fig. 2-G).
@@ III-D Commands
Commands define the goal of the task and need to be provided by the human operator or a high-level planner. Following previous work on pedipulation [5], we only command the foot position and allow the rest of the whole-body control to emerge during training. The foot position commands are sampled uniformly from the command space. In stage one, we apply a command space curriculum. In the subsequent stages, we provide separate command spaces for each obstacle scenario (Fig. 2).
While whole-body obstacle avoidance is desirable, we additionally need the option to allow the pedipulating foot to make contact with the environment to solve manipulation tasks, such as pushing doors open or moving objects. To this end, we provide an additional boolean command . During training, this switch determines whether penalties on the pedipulating foot’s contact forces are applied (Table III). The policy then learns to avoid obstacles when the switch is off and to allow contacts when it is on. Note that the remaining bodies of the robot always have contacts penalized and should thus always avoid obstacles.
@@ III-E Rewards
We apply this idea to all collision bodies of the robot to avoid exploitation strategies like pushing obstacles away with the hips, for example, which receive lower penalties than the pedipulating foot. Note that we choose a minimum force threshold for the contact events to avoid false positives due to numerical inaccuracies. We only penalize the non-pedipulating feet for contacts with obstacles but not the ground, as those forces should clearly be allowed. Finally, we give the contact rewards for the pedipulating foot based on the contact switch command (Sec. III-D).

### v3

@@ Abstract
Pedipulation leverages the feet of legged robots for mobile manipulation, eliminating the need for dedicated robotic arms. While previous works have showcased blind and task-specific pedipulation skills, they fail to account for static and dynamic obstacles in the environment. To address this limitation, we introduce a reinforcement learning-based approach to train a whole-body obstacle-aware policy that tracks foot position commands while simultaneously avoiding obstacles. Despite training the policy in only five different static scenarios in simulation, we show that it generalizes to unknown environments with different numbers and types of obstacles. We analyze the performance of our method through a set of simulation experiments and successfully deploy the learned policy on the ANYmal quadruped, demonstrating its capability to follow foot commands while navigating around static and dynamic obstacles. Videos of the experiments are available at sites.google.com/leggedrobotics.com/perceptive-pedipulation.
@@ I Introduction
Recent works have explored pedipulation – manipulation using a quadrupedal robot’s foot – to overcome these limitations [5, 6, 7]. These works use reinforcement learning (RL) to track position commands for the pedipulating foot with an emerging whole-body behavior. They successfully demonstrate skills such as pushing a button, opening a door, and lifting objects using the foot as an end-effector. However, these methods primarily rely on proprioceptive information and do not consider obstacles in the environment. Thus, the learned behaviors are unable to avoid unwanted collisions.
A potential approach to address this limitation is to develop an obstacle-avoiding high-level controller in a hierarchical architecture, leveraging the existing blind policies for low-level joint control. However, since the high-level planner can only command the pedipulating foot’s position without direct control over the remaining degrees of freedom (DoF), collisions between other parts of the robot and the environment may still occur during command tracking by the blind policy.
Thus, instead of a hierarchical design, this work proposes to extend the existing joint-level controllers by integrating perception in an end-to-end fashion. Such a policy retains the simple interface from prior works [5, 6, 7], requiring only a foot position command while extending the system’s capability to simultaneously perform whole-body obstacle avoidance. We train this policy entirely in simulation using simple arrangements of cuboidal obstacles and providing appropriate collision rewards to learn the desired collision avoidance behavior. In addition, we introduce a switch to allow or disallow contacts with the pedipulating foot. With this contact switch, the policy can either avoid obstacles entirely or perform pedipulation tasks while avoiding obstacles with the rest of the body, retaining the capabilities from our previous work [5].
@@ II Related Work
For general and precise manipulation tasks, recent works have looked into pedipulation - the manipulation of objects using a robot’s foot [6, 7, 5, 12]. These works specify a desired position for the pedipulating foot and control it with learning-based methods in a whole-body fashion. This approach provides a simple control interface for a human operator or a high-level planner to perform various tasks. However, the learned pedipulation policy is limited to obstacle-free environments, as it relies only on proprioceptive information. In this work, we aim to account for obstacles in the robot’s workspace by incorporating perceptive observations.
@@
Recently, researchers have also been investigating learning-based methods for obstacle avoidance [17]. Miki et al. [1] use perception to enable collision-free locomotion in confined spaces by training a high-level policy that commands the base pose and using a low-level policy to track this command. As motivated in Sec. I, this decoupling is limiting in the case of pedipulation. Honerkamp et al. [18] learn obstacle-avoiding manipulation skills for a wheeled-mobile base with an attached manipulator arm. In their work, the learned policy outputs the end-effector and base velocity commands, which are converted to the desired joint positions through inverse kinematics (IK). While it is possible to follow a similar approach that separates the system into a three-legged mobile base and the pedipulating leg, using IK reduces the whole-body reachability, as shown in [5].
@@ III-A Training Overview
We add Stage 3 to improve the sim-to-real transfer, as the elevation map on the robot is subject to noise, outliers, and occlusions. Training of a policy on all three stages takes roughly 6 days on an Nvidia RTX 4090.
@@ III-B1 Obstacles
Previous works targeting obstacle avoidance integrate obstacles into the static terrain [20, 21], which is computationally efficient as it avoids the computation of the obstacle dynamics. In the context of pedipulation, however, where moving obstacles should be possible when desired (Sec. III-D), training on static obstacles only would prevent the policy from performing interactive behaviors such as pushing or reorientation of obstacles during deployment. Thus, we use movable obstacles, which increase training times but provide representative experiences of collision dynamics.
@@ III-B2 Obstacle Scenarios
To encourage the policy to learn generalizable obstacle avoidance, we construct five obstacle scenarios (Fig. 2, Stage 2). We construct two scenarios using a single obstacle (Table LABEL:tab:obstacle_rand, Obstacle 1), focusing on close and far-range obstacle avoidance (Fig. 2-C and Fig. 2-D, respectively). Our third scenario (Fig. 2-E) has two obstacles (2 and 3) with a gap wide enough for the robot to spawn in between, which encourages the policy to act conservatively when constrained on both sides. Similarly, in the next environment (Fig. 2-F), we use two obstacles (2 and 3) to create a tight gap, into which the robot can only reach with its foot. Finally, using all three obstacles, we construct a scenario that features two gaps to reach through and position the obstacles such that there is enough space for the robot to turn on the spot to reach commands behind it (Fig. 2-G).
@@ III-D Commands
Commands define the goal of the task and need to be provided by the human operator or a high-level planner during deployment. Following previous work on pedipulation [5], we only command the foot position and allow the rest of the whole-body control to emerge during training. During training, the foot position commands are sampled uniformly from the current command space. In stage one, we apply a command space curriculum. In the subsequent stages, we provide separate command spaces for each obstacle scenario (Fig. 2). Over a single episode, we resample the commands three times with 6 seconds between each resample.
While whole-body obstacle avoidance is desirable, we additionally need the option to allow the pedipulating foot to make contact with the environment to solve manipulation tasks, such as pushing doors open or moving objects. To this end, we provide an additional boolean command . During training, this switch is set randomly for every episode and determines whether penalties on the pedipulating foot’s contact forces are applied (Table III). The policy then learns to avoid obstacles when the switch is off and to allow contacts when it is on. Note that the remaining bodies of the robot always have contacts penalized and should thus always avoid obstacles.
@@ III-E Rewards
We apply this idea to all collision bodies of the robot to avoid exploitation strategies like pushing obstacles away with the hips, for example, which receive lower penalties than the pedipulating foot. Note that we choose a minimum force threshold for the contact events to avoid false positives due to numerical inaccuracies. We only penalize the non-pedipulating feet for contacts with obstacles but not the ground, as those forces should clearly be allowed. Finally, we give the contact rewards for the pedipulating foot based on the contact switch command (Sec. III-D). All listed rewards were tuned manually by inspecting the performance of a trained policy and adjusting weights accordingly before retraining.